Jerry invested $1500 in an account that paid him 8.25% simple intrest what will the balance of his account be after 6 years

Answer by stanbon(75887)

Jerry invested $1500 in an account that paid him 8.25% simple intrest what will the balance of his account be after 6 years
---------
A(t) = P(1 + rt)
----
A(6) = 1500(1+0.0825*6) = 1500(1.495) = $2242.5
===================================================
Cheers,
Stan H.
=============